OpenStudy (anonymous):

I'm not a mathematics student but my friend need help in Maths f(x+1/x)= x-1/x so f(x)=?

OpenStudy (anonymous):

@mathslover @iambatman

Miracrown (miracrown):

let's assign g(x) = x - 1/x . h(x) = x + 1/x now, using function composition, f ( h(x) ) = g(x)